Grade 11Algebra

If 1/a+1/b+1/c=1/a+b+c then prove that 1/a^n+1/b^n+1/c^n= 1/a^n+b^n+c^n here n is any odd positive integer

Nishita
9 Years agoGrade 11

1 Answer

Harish 423
7 Years ago
\Bigl( \frac{1}{a} + \frac{1}{b}+ \frac{1}{c}\Bigr) (a+b+c) = 1\\ \\ \implies \sum \frac{1}{a}(b+c)+3=1 \\ \\ \implies \sum bc(b+c)+2abc=0 \\ \\ \implies (b+c)(c+a)(a+b)=0
 
Therefore one of (b+c),(c+a),or (a+c) is zero; so we can assume by symmetry , that b = -c
 
Now for any odd integers n , bn = – cn and therefore
 
\Bigl( \frac{1}{a^n} + \frac{1}{b^n}+ \frac{1}{c^n}\Bigr) = \frac{1}{a^n+b^n+c^n}